OpenStudy (anonymous):

Manuel scored 8 fewer points than Will scored. together they scored 44 points. How many points did Will score?

OpenStudy (anonymous):

44-8= 36 36/2=18 18+8= 26 Will scored 26 points. :)

OpenStudy (anonymous):

np :)